# Pentagram maps over rings, Grassmannians, and skewers

### v1

@@ 7.1 Closed polygons and super-periodic difference operators
A difference/pseudo-difference operator $\mathcal{D}=\sum\alpha_{i}T^{i}$ is said to be periodic, with period $n$, if all its coefficients $\alpha_{i}$ are $n$-periodic sequences. Equivalently, $\mathcal{D}$ is $n$-periodic if $\mathcal{D}T^{n}=T^{n}\mathcal{D}$. An $n$-periodic difference operator is said to be super-periodic if its kernel consists entirely of $n$-periodic sequences.
@@
If $\mathcal{D}$ is a super-periodic operator with period $n$, then $\varepsilon_{\mathcal{D}}$ is an $n$-periodic sequence in $(\mathrm{Ker}\,\mathcal{D})^{*}$. So, the mapping $\bar{\phi}\colon R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ takes super-periodic operator with period $n$ to closed polygons with the same period. Conversely, given a closed $n$-gon $(P_{i})$, we can choose spanning vectors $v_{i}\in P_{i}$ so that $v_{i+n}=v_{i}$, and then the operator $\mathcal{D}$ such that $\mathcal{D}v=0$ can be taken to be $n$-periodic. So, there is a surjective mapping the set of properly bounded super-periodic left difference operators of degree $d$ and period $n$ to the set of projective equivalence classes of closed $n$-gons . Furthermore, by Proposition 7.2 preimages of every class of polygons under this mapping are precisely the orbits of the left-right action of $n$-periodic sequences of units.∎
@@ 7.2 The trace of a periodic pseudo-difference operator
We need to show that $\mathrm{Tr}(\mathcal{Q}\widehat{\mathcal{Q}}-\widehat{\mathcal{Q}}\mathcal{Q})=0$. Note that the left-hand side is bilinear over $\mathbb{Z}$. Furthermore, if we endow the cyclic space with discrete topology, then the trace function becomes continuous, so the operation $\mathcal{Q},\widehat{\mathcal{Q}}\mapsto\mathrm{Tr}(\mathcal{Q}\widehat{\mathcal{Q}}-\widehat{\mathcal{Q}}\mathcal{Q})$ distributes over infinite sums. Hence, it suffices to verify the statement in the case $\mathcal{Q}=\alpha T^{k}$, $\widehat{\mathcal{Q}}=\beta T^{k}$, where $\alpha,\beta\in R^{\mathbb{Z}}$ are $n$-periodic. Clearly, $\mathrm{Tr}\,\mathcal{Q}\widehat{\mathcal{Q}}=\mathrm{Tr}\,\widehat{\mathcal{Q}}\mathcal{Q}=0$ if $l\neq-k$, so suppose $l=-k$. Then

### v2

@@ 7.1 Closed polygons and super-periodic difference operators
A difference/pseudo-difference operator $\mathcal{D}=\sum\alpha_{i}T^{i}$ is said to be periodic, with period $n$, if all its coefficients $\alpha_{i}$ are $n$-periodic sequences. Equivalently, $\mathcal{D}$ is $n$-periodic if $\mathcal{D}T^{n}=T^{n}\mathcal{D}$. An $n$-periodic difference operator is said to be super-periodic if its kernel consists entirely of $n$-periodic sequences, cf. [12].
@@
If $\mathcal{D}$ is a super-periodic operator with period $n$, then $\varepsilon_{\mathcal{D}}$ is an $n$-periodic sequence in $(\mathrm{Ker}\,\mathcal{D})^{*}$. So, the mapping $\bar{\phi}\colon R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ takes super-periodic operators with period $n$ to closed polygons with the same period. Conversely, given a closed $n$-gon $(P_{i})$, we can choose spanning vectors $v_{i}\in P_{i}$ so that $v_{i+n}=v_{i}$, and then the operator $\mathcal{D}$ such that $\mathcal{D}v=0$ can be taken to be $n$-periodic. So, the set of properly bounded super-periodic left difference operators of degree $d$ and period $n$ is mapped onto the set of projective equivalence classes of closed $n$-gons . Furthermore, by Proposition 7.2 preimages of every class of polygons under this mapping are precisely the orbits of the left-right action of $n$-periodic sequences of units.∎
@@ 7.2 The trace of a periodic pseudo-difference operator
We need to show that $\mathrm{Tr}(\mathcal{Q}\widehat{\mathcal{Q}}-\widehat{\mathcal{Q}}\mathcal{Q})=0$. Note that the left-hand side is bilinear over $\mathbb{Z}$. Furthermore, if we endow the cyclic space with discrete topology, then the trace function becomes continuous, so the operation $\mathcal{Q},\widehat{\mathcal{Q}}\mapsto\mathrm{Tr}(\mathcal{Q}\widehat{\mathcal{Q}}-\widehat{\mathcal{Q}}\mathcal{Q})$ distributes over infinite sums. Hence, it suffices to verify the statement in the case $\mathcal{Q}=\alpha T^{k}$, $\widehat{\mathcal{Q}}=\beta T^{{l}}$, where $\alpha,\beta\in R^{\mathbb{Z}}$ are $n$-periodic. Clearly, $\mathrm{Tr}\,\mathcal{Q}\widehat{\mathcal{Q}}=\mathrm{Tr}\,\widehat{\mathcal{Q}}\mathcal{Q}=0$ if $l\neq-k$, so suppose $l=-k$. Then
@@ 7.3 Invariants of the pentagram map
Remark 7.7
The ring of $n$-periodic pseudo-difference operators over $R$ is isomorphic to the ring $\text{M}_{n}(R)((z))$ of formal Laurent series over $\text{M}_{n}(R)$, cf. [8, Remark 3.8]. The isomorphism is given by
$\sum_{i=m}^{\infty}\alpha_{i}T^{i}\mapsto\sum_{i=m}^{\infty}\mathrm{diag}(\alpha_{i1},\dots,\alpha_{in})\Lambda^{i},$
where $\mathrm{diag}(\alpha_{i1},\dots,\alpha_{in})$ is the diagonal matrix with entries $\alpha_{i1},\dots,\alpha_{in}$, and
$\Lambda=zE_{n,1}+\sum_{i=1}^{n-1}E_{i,i+1}.$
Here $E_{i,j}\in\text{M}_{n}(R)$ is the matrix with a $1$ at position $(i,j)$ and zeros elsewhere. Therefore, in the periodic case, Lax representation (6.5) can be thought of as being valued in $\text{M}_{n}(R)((z))$. In other words, it is a Lax representation with spectral parameter. Also note that, for $R=\text{M}_{k}(\mathbb{F})$, we have $\text{M}_{n}(R)\simeq\text{M}_{nk}(\mathbb{F})$. So, in this case, one gets Lax matrices whose entries are in a field, allowing one to study the corresponding (Grassmann) pentagram maps using standard tools of integrable systems theory. This Lax representation is different from and in a sense dual to the one given in [4].